Ca(OH)2 + HCl → CaCl2 + H2O
How many grams of CaCl2 will be produced from 37 g of Ca(OH)2 and 37 g of HCl?
1
Expert's answer
2020-04-27T01:52:38-0400

The balansed reaction is as following:

Ca(OH)2 + 2HCl → CaCl2 + 2H2O

To find the mass of the product of the reaction, we need to determine a limiting reactant. For the purpose, we need to calculate the number of moles of Ca(OH)2 and HCl used in the reaction.

The number of moles equals:

n = m / Mr

where n - number of moles, m - mass, Mr - molecular weight.

From here:

n(Ca(OH)2) = 37 g / 74 g/mol = 0.5 mol

n(HCl) = 37 g / (2 × 36.5 g/mol) = 0.5 mol

The results show that there is no limiting reactant as the number of moles of both reactants are equal. As a result, the number of moles of CaCl2 produced in the reaction also equals 0.5 mol.

As n = m / Mr:

m(CaCl2) = n(CaCl2) × Mr(CaCl2) = 0.5 mol × 111 g/mol = 55.5 g


Answer: 55.5 g